# TEXasGAN: Tactile Texture Exploration and Synthesis System Using Generative Adversarial Network

### v2

@@ IV-A Data Preparation
Seven groups (G1–G4, G6, G8, and G9) were selected from the total of nine groups in LMT108. The G5 and G7 groups were excluded from the study, as the original stimuli were found to evoke a tactile sensation that was similar to that of other groups. And we selected two kinds of textures from each groups and built a training dataset containing fourteen classes.
@@
To transform spectrograms back into sound waves so that the vibrator can play them, we adopted Griffin-Lim algorithm [29] . The number of samples per frame was set at 2048, the iteration time was set to 50, the hop length was set to 2048 / 10 (floor rounding). To reduce the time cost and provide a smooth interaction experience, we employed the algorithm on GPU using Torchaudio [30].
@@
In order to train the model effectively, we segmented the spectrogram into fragments using a sliding window. This approach yielded several advantages, including dimensionality reduction and an augmentation in the training dataset’s size. The dimensions of the sliding window were set at 48 units along the frequency axis and 320 units along the time axis, corresponding to a segmentation of 48 $\times$ 320 spectrograms. The choice of 48 on the frequency axis corresponded to 1000 Hz, signifying that only the segment of the spectrogram within this frequency range was retained. Concurrently, the 320 units along the time axis approximated a duration of 1.5 seconds for each spectrogram segment. This temporal length was deemed sufficient to capture the full information of the underlying audio signal.
@@
The original dataset contains about five seconds sound recordings of different surfaces. Here we used a sliding window of 48 $\times$ 320 corresponding to about 1.5 seconds, and moved it on transformed spectrograms, so that we can obtain segmentations from different beginning time. The moving step length was set at five units (about 23 milliseconds), and produced a total of 24480 spectrogram segments.
@@
IV-B Model Training
The size of the latent vector $z$ was set to 128. In our attempts, larger latent vector cannot improve the performance obviously, but caused fluctuations and a longer convergence time. Network training hyperparameters were defined as follows: a batch size of 128, the learning rates of $G$ and $D_{s}$ were set to 0.0002, and the $E$, $C$ and $D_{z}$ were set to 0.001. The learning rates of $C$ and $D_{z}$ was decreased by a factor of 0.95 with the increasing of the epoch number. We opted for binary cross-entropy for the adversarial loss and cross-entropy for the auxiliary classification loss, while for the decoder loss, Mean Square Error (MSE) was chosen. The Adam optimizer was employed for efficient optimization. In order to mitigate overfitting, we utilized soft labels within the adversarial component. Specifically, a one-sided label smoothing was implemented, with a soft scale parameter set to 0.3. This signifies that random values between 0.7 and 1 were utilized instead of a fixed value of 1, and similarly, random values between 0 and 0.3 were employed instead of 0. The entire training regimen spanned 50 epochs. After the training process, only the generator part was used the optimization system. The training process was conducted on an RTX A6000 GPU, took 2.5 hours.
In order to expedite convergence, mitigate bias, and enhance stability, we applied min-max normalization to the training dataset, subsequently rescaling the range of values to fall within the [-1, 1] range. This normalization process aligns the data range of both the input and output of the generator, facilitating the training process. Additionally, we performed the inverse transformation to revert the network’s output range back to the original real data range, ensuring that the generated data remains consistent with the original data distribution.
IV-C Initialization
A better initial point will help the optimizer find the target well. DSS has attempted initialization approaches such as providing options or limiting the distance between initial point and target to control the difficulty of the task and avoid bad initialization [9].
There is a more severe challenge in our optimization setup. While the auxiliary classifier enhances class distinguishability, it widens the gap between classes, potentially hindering the optimizer from traversing between clusters and restricting it to a confined space (details of the latent space distribution in Sec. V-A).
@@ IV-C Initialization
To bring $z$ closer to the target area, the principle of our search strategy is: Choose near but different new $z$ for “So-so”, and far new $z$ for “Bad”. The average distance of vectors in the latent space $dis_{avg}$ was calculated. Here we define a parameter $step=\frac{1}{8}\times dis_{avg}$ as the unit of the moving distance during the exploration. Inspired by the tabu search, visited latent vectors are added to the tabu list. In this study, a specified radius $r$ is enforced for vectors within the tabu list (here $r$ is defined as the average distance $step$). Consequently, vectors within this range are prohibited, allowing for the exploration of larger regions. In this study, the length of the tabu list is set to five, meaning that the areas surrounding the five most recently visited latent vectors are prohibited from the further exploration. Following the search principle mentioned, when the user selects “So-so”, we choose a new $z$ within the range of (0.5 $\times$ $step$, 2 $\times$ $step$) from the current $z$. If the user chooses “Bad”, we select a new $z$ from distances beyond 2 $\times$ $step$ from the current $z$.
This simple rule is easy to implement and allows for rapid execution. Through this method, we can efficiently select an initial latent vector that is closer to the target region, mitigating the risk of not being able to reach the target area due to gaps in the distribution of classes.
Figure 7: The distribution visualization of the latent space.
IV-D User Interface and Presenting Device
Fig. 4 and Fig. 5 show the vibrotactile display for evaluating the designed vibration and the user interface of DSS. The device [12] was composed of a vibro-transducer (Acouve Vp210), a power amplifier (BOSE 1705II), sponge and an acrylic cover board. The vibrator was set in the sponge block so that the vibration can be transmitted effectively. The acrylic board covered with tapes was pasted onto the top surface of the vibrator, so that users can place their fingers on the plane surface and feel the generated vibration.
During the optimization process, users can click the “Play” button above to play the actual vibration recording serving as the target. When users drag the slider below, vibrations generated from the feature vector corresponding to the slider’s position will be played. Users can freely explore the high-dimensional feature space by moving the one-dimensional slider. Once users locate the slider position that best matches the target vibration, they can click “Next” to proceed to the next iteration. Users can also click “Save” at any time to save the currently generated vibration.
Figure 8: The figure shows one of the target vibrations for each class, along with the corresponding vibrations generated by 3 users.
IV-E Stage I: Generation
To assess the generative performance of our TEXasGAN, we designed a two-stage experiment that includes a generation experiment and an evaluation experiment. To validate whether TEXasGAN can generate new samples different from the training data, we selected five classes from LMT108 that differ from the training set as generation targets, as shown in Fig. 6.
Due to the free movement of the collection device on surfaces during the vibration acquisition process [28], vibration is not uniformly distributed on the time axis. This implies that not all samples can effectively represent the characteristics of a particular class. To provide subjects with more reliable target vibrations, facilitating better execution of the generation task, we manually selected three relatively uniform and representative vibrations for each class. This resulted in a total of fifteen target vibrations, and subjects were tasked with generating samples as similar as possible to these vibrations.
@@ IV-E Stage I: Generation
After subjects familiarized themselves with the system and the characteristics of vibrations, we proceeded to the generation experiment process. Subjects were required to generate a sample corresponding to each target vibration. The first step was the initialization of the optimization process. In this step, subjects needed to use the options “Bad”, “So-so” and “Good” to select a suitable initial value for the optimizer. During this process, subjects were only required to make rough comparisons, so they were instructed not to focus on detailed perceptions, but to perform quick and multiple selections, with the aim of covering as broad an exploration range as possible through multiple attempts. After entering the optimization process, subjects were still encouraged to make as many selections as possible rather than sticking to a single iteration’s slider position selection. However, in this phase, subjects were asked to carefully perceive various aspects of vibrations, such as frequency and amplitude. Moreover, to distinguish between different slider positions, this optimization process might be more time-consuming. When consecutive iterations did not produce vibrations significantly different along the slider direction, subjects could adopt several strategies:
@@
    If the subject felt dissatisfied with the current vibration, restart the optimization process.
Following strategy 1 or strategy 2 for several consecutive iterations usually allowed the optimizer to move away from the current state. During the experiment process, subjects were asked to freely choose strategies based on their subjective judgment. The entire process of the generation experiment lasted approximately 2 hours, including rest breaks. Each subject generated a total of 15 samples corresponding to the target vibrations. From initialization to completion, the generation of each sample took 2-8 minutes.
In order to cover as many generated results as possible in the subsequent pairing task of the evaluation experiment without overwhelming the subjects, we limited the number of generated samples. In the first stage, we invited three subjects (three males, aged 21-26), who totally generated forty-five samples (nine for each class). These three subjects will not be involved in the evaluation phase to eliminate the impact of memory.
@@ IV-F Stage II: Evaluation
1. Select the correct class for the real vibration (the control group): In this part, all fifteen real vibrations were presented in a shuffled order to the subject. For each trial, one real vibration was selected for classification, and subjects could click a button to play this vibration. Also, five real vibrations from each class were presented. These were randomly selected from each class, excluding the vibration to be classified shown above. The indices (1, 2, 3, 4, 5) corresponding to the vibration classes (G2, G3, G4, G6, G8). Subjects could play these vibrations and were required to choose the class that most closely resembled the one being classified. This process was repeated three times, resulting in a total of forty-five classifications. This part aimed to validate the study’s premise that people can perceive differences between vibration stimuli, understand their characteristics, and accurately classify them. The results from this control group were used to assess the accuracy of the generated vibrations. Subjects were also asked to describe the features of the five types of vibrations and perform subsequent tasks based on these descriptions.
2. Select the correct class for the generated vibration. The procedure for this part was similar to the first, but focused on classifying generated vibrations. The forty-five generated vibrations were shuffled and presented to the subjects for classification. The vibrations below were still randomly selected from the real vibrations of each class. Subjects were required to select the class that most closely resembled the vibration generated from these five options. This process was only conducted once to maintain an equal task number with the control group. We asked subjects to describe the vibration features and compare with the control group.
3. Select the corresponding target vibration for the vibration generated. The third part aimed to validate whether subjects can identify differences between samples in the same class. The forty-five shuffled generated vibrations were shown to the subjects, same as in the second part above. However, instead of five options, subjects were presented with three options, all belonging to the same class as the displayed generated vibration. The reason is, for each class, there were three real vibrations serving as targets. Therefore, subjects needed to select the correct target vibration from the below three options corresponding to the given generated vibration. The total number of classification tasks in this part was also forty-five, decided by the number of generated vibrations.

### v4

@@ IV-A Data Preparation
Seven groups (G1–G4, G6, G8, and G9) were selected from the total of nine groups in LMT108. The G5 and G7 groups were excluded from the study, as the original stimuli were found to evoke a tactile sensation that was similar to that of other groups [13]. While more data samples could enhance the generative model’s learning, highly complex data can also hinder model convergence. Additionally, since we used an ACGAN-like auxiliary classifier to control the generation process, using the full LMT108 dataset would require the classifier to learn a challenging 108-class classification. Although CGAN could incorporate conditional information without the need for an auxiliary classifier, after comparing its generative performance (discussed further in Sec. V-B), we finally chose to train the model using a subset of the dataset’s categories. We selected two representative types of textures from each group and built a training dataset containing fourteen classes, to ensure the model could learn a sufficiently diverse range of features, thus enhancing its capacity for generalized knowledge. The selected classes are shown in Table I.
TABLE I: Selected classes of training set.
Group Class
G1 Rhomb Aluminum Mesh, Squared Aluminum Mesh
G2 Granite Type Veneziano, Roof Tile
G3 Aluminum Plate, Ceramic Plate
G4 Bamboo, Laminated Wood
G6 Carpet, Fiber
G8 Cardboard, Wall Paper
G9 Floor Cloth, Kashimir
@@
To transform spectrograms back into sound waves to ensure that the vibrator can play them, we adopted Griffin-Lim algorithm [31]. The number of samples per frame was set at 2048; the iteration time was set to 50; the hop length was set to 2048 / 10 (floor rounding). To reduce the time cost and provide a smooth interaction experience, we employed the algorithm on GPU using Torchaudio [32].
@@
To train the model effectively, we segmented the spectrogram into fragments using a sliding window. This approach yielded several advantages, including dimensionality reduction and an augmentation in the training dataset’s size. The dimensions of the sliding window were set at 48 units along the frequency axis and 320 units along the time axis, corresponding to a segmentation of 48 $\times$ 320 spectrograms. The choice of 48 on the frequency axis corresponded to 1000 Hz, signifying that only the segment of the spectrogram within this frequency range was retained. Concurrently, the 320 units along the time axis approximated a duration of 1.5 seconds for each spectrogram segment. This temporal length was deemed sufficient to capture the full information of the underlying audio signal.
@@
The original dataset contains approximately 5 seconds sound recordings of different surfaces. Here, we used a sliding window of 48 $\times$ 320 corresponding to approximately 1.5 seconds, and moved it on transformed spectrograms, to ensure that we can obtain segmentations from different beginning time. The moving step length was set at five units (about 23 milliseconds), and produced a total of 24480 spectrogram segments.
IV-B Model Training
The size of the latent vector $z$ was set to 128. In our attempts, larger latent vector cannot improve the performance, but caused fluctuations and a longer convergence time. Network training hyperparameters were defined as follows: a batch size of 128, the learning rates of $G$ and $D_{s}$ were set to 0.0002, and the $E$, $C$ and $D_{z}$ were set to 0.001. The learning rates of $C$ and $D_{z}$ was decreased by a factor of 0.95 with the increasing of the epoch number. We opted for binary cross-entropy for the adversarial loss and cross-entropy for the auxiliary classification loss, while for the decoder loss, Mean Square Error (MSE) was chosen. The Adam optimizer was employed for efficient optimization. To mitigate overfitting, we utilized soft labels within the adversarial component. Specifically, a one-sided label smoothing was implemented, with a soft scale parameter set to 0.3. This signifies that random values between 0.7 and 1 were utilized instead of a fixed value of 1, and similarly, random values between 0 and 0.3 were employed instead of 0. The entire training regimen spanned 50 epochs. After the training process, only the generator part was used for the optimization system. The training process was conducted on an RTX A6000 GPU, and took 2.5 hours.
To expedite convergence, mitigate bias, and enhance stability, we applied min-max normalization to the training dataset, subsequently rescaling the range of values to fall within the range of [-1, 1]. This normalization process aligns the data range of both the input and output of the generator, facilitating the training process. Additionally, we performed the inverse transformation to revert the network’s output range back to the original real data range, ensuring that the generated data remain consistent with the original data distribution.
IV-C Initialization
An ideal initial point will help the optimizer find the target well. DSS has attempted initialization approaches, such as providing options or limiting the distance between initial point and target to control the difficulty of the task and avoid bad initialization [10].
There is a critical challenge in our optimization setup. While the auxiliary classifier enhances class distinguishability, it widens the gap between classes, potentially hindering the optimizer from traversing between clusters and restricting it to a confined space (details of the latent space distribution in Sec. V-A).
To bring $z$ closer to the target area, the principle of our search strategy is: Choose near but different new $z$ for “So-so”, and far new $z$ for “Bad”. The average distance of vectors corresponding to the training set samples in the known latent space $dis_{avg}$ was calculated. Here, we define a parameter $step=\frac{1}{8}\times dis_{avg}$ as the unit of the moving distance during the exploration. Inspired by the tabu search, visited latent vectors are added to the tabu list. In this study, a specified radius $r$ is enforced for vectors within the tabu list (here $r$ is defined as the average distance $step$). Consequently, vectors within this range are prohibited, allowing for the exploration of larger regions. In this study, the length of the tabu list is set to five, meaning that the areas surrounding the five most recently visited latent vectors are prohibited from the further exploration. Following the search principle mentioned, when the user selects “So-so”, we choose a new $z$ within the range of (0.5 $\times$ $step$, 2 $\times$ $step$) from the current $z$. If the user chooses “Bad”, we select a new $z$ from distances beyond 2 $\times$ $step$ from the current $z$.
This simple rule is easy to implement and allows for rapid execution. Through this method, we can efficiently select an initial latent vector that is closer to the target region, mitigating the risk of not reaching the target area owing to gaps in the distribution of classes.
IV-D User Interface and Presenting Device
Fig. 4 and Fig. 5 show the vibrotactile display for evaluating the designed vibration and the user interface of DSS. The device [13] was composed of a vibro-transducer (Acouve Vp210), a power amplifier (BOSE 1705II), sponge and an acrylic cover board. The vibrator was set in the sponge block to ensure that the vibration can be transmitted effectively. An acrylic plate was attached to the vibrator, with tape covering the upper surface of the acrylic to increase friction. This setup allowed participants’ fingers to make closer contact with the vibrating acrylic plate, helped them perceive the vibrations.
@@ IV-D User Interface and Presenting Device
During the optimization process, users can click the “Play” button above to play the actual vibration recording serving as the target. When users drag the slider below, vibrations generated from the feature vector corresponding to the slider’s position will be played. Users can freely explore the high-dimensional feature space by moving the one-dimensional slider. When users locate the slider position that best matches the target vibration, they can click “Next” to proceed to the next iteration. Users can also click “Save” at any time to save the currently generated vibration.
IV-E Stage I: Generation
To assess the generative performance of our TexSenseGAN, we designed a two-stage experiment that includes a generation experiment and an evaluation experiment. To validate whether TexSenseGAN can generate new samples different from the training data, we selected classes differ from the training set as generation targets. Although we following previous research used 7 groups in the training process, our further observed that G1 and G9 exhibited excessively weak vibration intensities or significant similarity to other categories from the subjective feeling. Consequently, we also excluded G1 and G9. Finally, following stratified sampling principles, we randomly selected one class from each of the remaining groups for the task. The selected classes are shown in Fig. 6.
Owing to the free movement of the collection device on surfaces during the vibration acquisition process [30], vibration is not uniformly distributed on the time axis. Therefore, not all samples can effectively represent the characteristics of a particular class. To provide subjects with more reliable target vibrations, facilitating better execution of the generation task, we manually selected three relatively uniform and representative vibrations for each class. This resulted in a total of fifteen target vibrations, and subjects were tasked with generating similar samples to these vibrations.
@@ IV-E Stage I: Generation
After subjects familiarized themselves with the system and characteristics of vibrations, we proceeded to the generation experiment process. Subjects were required to generate a sample corresponding to each target vibration. The first step was the initialization of the optimization process. In this step, subjects needed to use the options “Bad”, “So-so”, and “Good” to select a suitable initial value for the optimizer. During this process, subjects were only required to make rough comparisons; thus, they were instructed not to focus on detailed perceptions, but to perform quick and multiple selections, with the aim of covering an exploration range through multiple attempts. It is important to note that the three options here do not represent the actual quality of the samples. For instance, “Good” does not mean that the sample is sufficiently close to the target sample, but rather that it possesses certain characteristics similar to the target (e.g., similar frequency components), making it a suitable initial point for optimization. After entering the optimization process, subjects were encouraged to make numerous selections rather than sticking to a single iteration’s slider position selection. However, in this phase, subjects were asked to carefully perceive various aspects of vibrations, such as frequency and amplitude. We introduced several effective ways to perceive signal changes and adjustment strategies to help participants better understand the generation process and complete the task more easily. Like participants can focus on adjusting one parameter at a time, for example, the frequency, and then other aspects. This makes the changes during the adjustment process more noticeable. Participants are also advised to carefully adjust the sliders to avoid generating distorted vibration signals. Distorted signals typically indicate that the exploration has exceeded the model’s acceptable latent space, which can lead to generation failure. Therefore, such outcomes should be avoided as much as possible. Moreover, to distinguish between different slider positions, this optimization process might be more time-consuming. When consecutive iterations did not produce vibrations significantly different along the slider direction, subjects could adopt several strategies:
@@
    If the subject felt dissatisfied with the current vibration, restart the optimization process from the initialized point.
Following strategy 1 or strategy 2 for several consecutive iterations usually allowed the optimizer to move away from the current state. During the experiment process, subjects were asked to freely choose strategies based on their subjective judgment. However, we encourage participants to primarily opt for strategies 1 or 2 and minimize the use of the reset operation during the optimization process. If participants feel fatigued, prefer not to continue with the iterative adjustments of strategies 1 or 2, or if the generated samples become distorted, they may then choose strategy 3. Since carefully perceiving subtle differences in vibration can be mentally exhausting, and prolonged exposure may cause finger numbness, potentially affecting sensitivity, participants are allowed to take breaks at any point during the experiment to maintain accuracy in their perceptions. In our experimental observations, the reset operation was used no more than three times across all user trials.
During the selection of initial points and the optimization of vibrations, participants were also asked to rate the subjective similarity between the sample and the target at three different time points: the initial random sample, the sample after initialization, and the optimized sample. Ratings were given on a scale of 0 to 10 (0 = completely different, 10 = highly similar). Unlike the options in the initialization mentioned earlier, participants here were required to directly assess the degree of similarity between the current sample and the target. The entire process of the generation experiment lasted approximately 2 hours, including rest breaks. Each subject generated 15 samples corresponding to the target vibrations. From initialization to completion, the generation of each sample took 2-8 minutes.
To cover multiple generated results in the subsequent pairing task of the evaluation experiment without overwhelming the subjects, we limited the number of generated samples. In the first stage, we invited three subjects (three males, aged 21-26), who generated 45 samples (nine for each class). These three subjects will not be involved in the evaluation phase to eliminate the impact of memory.
@@ IV-F Stage II: Evaluation
Figure 8: The distribution visualization of the latent space.
Figure 9: The figure shows the subjective similarity of each vibration class in different optimization stages.
1. Select the correct class for the real vibration (the control group): In this part, all 15 real vibrations were presented in a shuffled order to the subject. For each trial, 1 real vibration was selected for classification, and subjects could click a button to play this vibration. Furthermore, 5 real vibrations from each class were presented. These were randomly selected from each class, excluding the vibration to be classified shown above. The indices (1, 2, 3, 4, 5) corresponding to the vibration classes (G2, G3, G4, G6, G8). Subjects could play these vibrations and were required to choose the class that most closely resembled the one being classified. This process was repeated three times, resulting in a total of 45 classifications. This part aimed to validate the study’s premise that people can perceive differences between vibration stimuli, understand their characteristics, and accurately classify them. The results from this control group were used to assess the accuracy of the generated vibrations. Subjects were asked to describe the features of the five types of vibrations and perform subsequent tasks based on these descriptions.
2. Select the correct class for the generated vibration. The procedure for this part was similar to the first, but focused on classifying generated vibrations. The forty-five generated vibrations were shuffled and presented to the subjects for classification. The vibrations below were randomly selected from the real vibrations of each class. Subjects were required to select the class that most closely resembled the vibration generated from these five options. This process was only conducted once to maintain an equal task number with the control group. We asked subjects to describe the vibration features and compare with the control group.
3. Select the corresponding target vibration for the vibration generated. The third part aimed to validate whether subjects can identify differences between samples in the same class. The forty-five shuffled generated vibrations were shown to the subjects, same as in the second part above. However, instead of five options, subjects were presented with three options, all belonging to the same class as the displayed generated vibration. This is because, for each class, there were three real vibrations serving as targets. Therefore, subjects needed to select the correct target vibration from below three options corresponding to the given generated vibration. The total number of classification tasks in this part was also 45, determined by the number of vibrations generated.
@@ V-A Encoded Latent Space
V-B Problems in Other Alternatives
In addition to the network model used in TexSenseGAN, we implemented various generators or decoders for comparison, including basic GAN, autoencoder; Variational Autoencoder (VAE) to learn the distribution; AAE to align the distribution, CGAN, and ACGAN to incorporate conditional information. However, these models faced challenges operating within our optimization framework, making direct comparisons unfeasible here.
Initially, we experimented with models based on the autoencoder architecture, including autoencoders, VAEs, and AAEs. While these models were relatively easy to train, they suffered from severe over-smoothing issues. Considering our need for feature vectors suitable for optimization, we targeted a relatively small feature space dimension, such as the 128 dimensions used in this study. However, this dimensionality proved inadequate for the original data, resulting in significant information loss when encoded into this latent space. Consequently, the generated spectrograms lacked detailed information, leading to noticeable differences between the reconstructed vibrations and the originals.
Figure 10: The figure shows one of the target spectrogram for each class, along with the corresponding spectrograms generated by 3 users.
Figure 11: The figure shows the waveform of the vibrations corresponding to the spectrograms in Fig. 10
Moreover, we incorporated GANs into our framework. Although GANs could generate spectrograms with more detail compared to autoencoders, they were prone to mode collapse, producing similar results for different inputs. Thus, GANs were unsuitable for our optimization approach based on input vectors. Additionally, GANs generate results from random noise, resulting in a mixed latent space. Consequently, the generated vibrations underwent significant changes when users moved the slider continuously. Contrary to the original DSS research, which focused on visual stimuli, our work emphasized on tactile tasks, which were less sensitive than visual stimuli. Consequently, these abrupt changes made it difficult for users to touch and compare vibrations, complicating the task of selecting an appropriate slider position.
We compared generative models, CGAN and ACGAN, with conditional information support. Within our task, unlike ACGAN, CGAN only requires conditional information as input, thus avoiding the challenges of multi-class classification and allowing it to handle larger datasets. However, our tests demonstrated that this approach, where conditional information is used alongside latent vectors to control generation, was less effective than ACGAN’s direct use of an auxiliary classifier to shape the latent space distribution. Under our training framework, CGAN tended to produce extremely sparse output results or exhibited a tendency not to converge. ACGAN performed relatively better among the control group models we experimented with; however, it faced challenges when applied in the optimization process, such as DSS. ACGAN can differentiate features between different classes and control them, but the uneven input vector format of label + latent vector is not conducive to the optimization process of DSS. Despite these challenges with ACGAN, we incorporated a similar auxiliary classification approach in TexSenseGAN, leveraging its strong capability for latent space control.
These problems make it difficult to identify models with similar performance for comparison with our TexSenseGAN. Therefore, we excluded the performance results or user studies of the control group.
V-C User Study
The Fig. 9 illustrates changes in subjective similarity to the target sample across three stages of the generation process: “Random” represents the initial random latent vector, “Initialized” denotes samples that have been initialized with certain target characteristics, and “Optimized” indicates the final optimized result. Light thin lines show changes in each individual trial (9 trials per class, with 3 participants generating samples for 3 target samples). The dark thick line represents the average subjective similarity for each class, with error bands indicated. Overall, the optimization process demonstrates effectiveness, supported by the benefits of the initial value selection method. Nevertheless, certain cases reveal instances of ineffective or even inverse optimization, with variations in effectiveness across different classes. Also, some generated spectrogram samples of Stage I are shown in Fig. 10, and the corresponding waveforms transformed from the spectrograms are shown in Fig. 11, to show the generation performance.